OpenStudy (anonymous):

Prove: 32(sin^4x)(cos^2x)=2-cos2x-2cos4x+cos6x Please help :(

sam (.sam.):

The RHS has all the cosine function, you might wanna change the LHS sine function to cosine

sam (.sam.):

\[32(\sin^4x)(\cos^2x)\] \[\Rightarrow32(\sin^2(x)\sin^2(x))(\cos^2x)\] Using \[\sin^2x=1-\cos^2x\] \[\Rightarrow32((1-\cos^2x)(1-\cos^2x))(\cos^2x)\]

sam (.sam.):

Then just distribute them you might get RHS

OpenStudy (loser66):

@lesliesaree can you get it? I cannot. my answer is not the same your right hand side although I apply .Sam 's instruction

OpenStudy (loser66):

\[32(1-cos^2)(1-cos^2)(cos^2)= 32(1-cos^2)^2(cos^2)\] \[= 32(1-2cos^2 +cos^4)(cos^2)= 32(cos^2-2cos^4+cos^6)\] \[= 32cos^2 -64cos^4+32cos^6\] it's not the same your RHS

OpenStudy (anonymous):

No I am not getting it because it needs to be reduced from the powers rather then ending with huge exponents but I keep getting stuck

OpenStudy (loser66):

your RHS is \[cos^2 x ~or~ cos (2x)\]

OpenStudy (loser66):

something else, just example

OpenStudy (loser66):

why didn't you ask him when you didn't get?

sam (.sam.):

If RHS is \[2-\cos(2x)-2\cos(4x)+\cos(6x)\] Prove RHS = LHS is easier and Then you'll have to use some of these \[\Rightarrow \cos(2x)=1-2\sin^2(x) \\ \\ \Rightarrow \cos(2x)=2\cos^2(x)-1\] And \[\cos(A+B)=\cos(A)\cos(B)-\sin(A)\sin(B)\]

OpenStudy (loser66):

I am ok with this stuff, thanks Sam

OpenStudy (anonymous):

Sorry :( but I was going to try that but I don't know how to reduce the ones that are bigger then 2 :o